give the dimensions of rectangles with a perimeter of 70 feet and length-to-width ratios of 3to4 4to5 and 1to1
Answer by rfer(10417)

P=2w+2l
p=2(3x)+2(4x)=70
6x+8x=70
14x=70
x=5
3x=15
4x=20
--------------
2(4x)+2(5x)=70
8x+10x+70
18x=70
x=3.89
4x=15.6
5x=19.4
----------------
2x+2x=70
4x=70
x=17.5
4x=70